OpenStudy (anonymous):

Glenn works as an office clerk and his job pays him $12 per hour. Shelia works as a cashier and her job pays her $13.50 per hour. Together, they worked 55 hours and earned a total of $705. How many hours did Glenn work in his job? A. 20 B. 30 C.35 D.25

OpenStudy (zpupster):

Shelia=x hours Glenn=y hours we know x+y = 55 13.50 x and 12y taking the number of hours times amount per hour will equal 705 so are system of equations are x+y = 55 13.50x + 12y = 705

OpenStudy (anonymous):

Can you explain that to me a bit more please

OpenStudy (zpupster):

ok you probably are having trouble solving the system of equations right??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

My problem is, I dont understand how the problem is supposed to be set up so I can figure out the question

OpenStudy (zpupster):

ok multiply x+y=55 times -12 -12(x+y=55) -12x -12y = -660 and add to the next equation 13.50x + 12y = 705 ____________________________ 1.50x = 45 divide both sides by 1.50 x= 30 this is sheila's hours worked what is Glen's

OpenStudy (zpupster):

do you understand all the steps

OpenStudy (anonymous):

No

OpenStudy (zpupster):

ok i looked at the total number of hours worked 55 and i assigned a x to sheila and y to glen and then formed an equation x+y = 55 so far so good

OpenStudy (anonymous):

Alright I'm understanding that part now

OpenStudy (zpupster):

all right next if you take the number of hours a person worked and multiply it by the amount they get paid per hour that is their total amount paid. but here we have two people so we add their amounts together remember sheila is x and glen is y so we have to multiply their no hours worked by what there hourly rate is sheila 13.50 * x glen 12 * y this will equal the total amount 705 dollars here is the second equation 13.50x + 12 y = 705 dollars

OpenStudy (zpupster):

ok with that??

OpenStudy (anonymous):

Yes!!! Thank you very much!